OpenStudy (anonymous):

how do you find the side lengths of a 30 60 90 triangle?

OpenStudy (anonymous):

By using \(\text{Pythagorean Theorem..}\)

OpenStudy (anonymous):

Or more precisely, with pen and paper and hands too.. :

OpenStudy (anonymous):

sin(30 degrees)=1/2 Hypotenuse is 2, the side opposite the 30 degree angle is 1 and the remaining side is\[\sqrt{3} \]

OpenStudy (anonymous):

Where is xx109 ?

OpenStudy (gorv):

use one angle and find ratio of any two side

OpenStudy (gorv):



OpenStudy (gorv):

sin30=1/2=perpendicular/hypotenuse=a/b

OpenStudy (gorv):

cos30=base/hypotenuse=a/c

OpenStudy (gorv):

\[\frac{ 1 }{ 2 }=\frac{ a }{ b }\] \[b=2a\]

OpenStudy (gorv):

\[c=\frac{ a }{ \frac{ \sqrt3 }{ 2 }}\]

OpenStudy (gorv):

now use cosine formula

OpenStudy (gorv):

u know cosine formula

OpenStudy (gorv):